a container containing a liquid is sliding down a inclined plane determine the angle made by the surface of the liquid with the horizontal during the motion

Grade:12

4 Answers

Nishant Vora IIT Patna
7 years ago
tan\theta = -g/a

This is the formula i guess for finding angle

Thanks
GAGANDEEP SINGH
7 years ago
The angle made by the liquid with plane is given by tan x = g/a, here a is acceleration of the liquid and g is acceleration due to gravity. Now, here rotate your frame by angle y of the plane so that g become gcos y and a becomes g siny so angle would be complement of y.
Ankit Jaiswal
7 years ago
but tell me how to prove it
GAGANDEEP SINGH
7 years ago
Open Page Number 262 of Concepts Of Physics Volume 1, By H C Verma Sir. There the given formula is proved.